# references.bib

### v3

@@ 5.1 Boolean algebras
Our general method in handling subvarieties will be to take the Vietoris complex construction, and show that by quotienting by additional equations, one can obtain sub-complexes which limit to the free algebra in the desired variety. To illustrate this, we consider the simple case of Boolean algebras: let $\mathcal{D}$ be a distributive lattice. Consider $\mathbf{F}_{DLat}(\{(a,b):a,b\in D\})/\Theta\cup\{p\vee\neg p=1\}$, the free algebra quotiented as before, with the addition of the equation $p\vee\neg p$, with which one forces Booleannness. Dually, this takes a given subspace of the space $V_{1}(X)$:
Lemma 38.
Let $X$ be a Priestley space. Given a point $C\in V_{1}(X)$, we have that $C\in[\neg U]\cup[U]$ for each $U$ if and only if $C$ is a singleton.
@@
Using Lemma 38 one can prove the following:
Theorem 39.
For each Priestley space $X$, $B_{1}(X)\cong X$ as Stone spaces, and $B_{1}(X)$ is isomorphic to the free Stone space generated by $X$.
@@
The isomorphism is clear; the freeness follows by the same arguments as in Lemma 20. ∎
@@
It is also worthy to note that Boolean algebras have a peculiar feature as far as our construction $V_{G}$ goes: if $X$ is a discrete Priestley space (definitionally equivalent to a Stone space), then $V_{G}(X)\cong X$, and in fact $V_{1}(X)\cong X$ already (since rooted subsets of singletons have to simply be singletons as well). Stone spaces also exhibit some further interesting properties in this respect:
Theorem 40.
@@
    $V_{1}(X)\cong X$;
@@
Assume that $X$ is not a Stone space, i.e., there is a pair of points $x,y$ such that $x\leq y$. Note that then $x<y$ is a subposet of $X$. Recall from Example 10 that from such a poset one constructs the whole Rieger-Nishimura ladder, and there is throughout each layer of the construction a point $z_{n}$ such that $r$ does not provide an isomorphism. Since all sets involved are finite, this means that $V_{n}(X)$ will always likewise contain such a point $z_{n}$ where $r$ cannot provide an isomorphism.
@@ 5.2 KC-algebras
In order to generalise our construction to $\mathsf{KC}$, we will need first the notion of a $\mathsf{KC}$-distributive lattice. A dual perspective can illustrate what these should be: they should be the Priestley spaces which sit on a directed poset, precisely the distributive lattices over which the structure of a KC-algebra can be built; as it happens, this admits a simple dual description:
Definition 41.
Let $(X,\leq)$ be a Priestley space. Let $R=(\leq\cup\leq^{\frown})^{*}$.We say that a subset $C$ is order-connected if for each $x,y\in C$, we have $xRy$. We say that $C$ is an order-component if it is a maximal order-connected subset.
Proposition 42.
Let $X$ be a Priestley space. Then the following are equivalent:
  1. 1.
    $X$ is directed as a poset;
  2. 2.
    If $C\subseteq X$ is an order-component, then $C$ has a unique maximal element.
@@
It is clear, since each poset can be decomposed into its order-components, that if each order-component has a unique maximal element, then it is directed. Now assume that $(X,\leq)$ is directed. First, note that order-components are closed subsets: if $x\notin C$, then pick a clopen upset containing $x$, and this will provide an open neighbourhood separate from $C$. So by the theory of Priestley spaces, $\max(C)\neq\emptyset$. Now if $x,y\in\max(C)$, and they are different, since $C$ is an order component, then there must be some point $z$ such that $z\leq x$ and $z\leq y$ – which contradicts directedness. Hence we have the result. ∎
We will use the above lemma freely, and refer to Priestley spaces in these conditions simply as “directed”.
Let us first outline our construction in algebraic terms. Given an appropriate distributive lattice $\mathcal{D}$, consider the sequence:
$(\mathcal{D},H(\mathcal{D}),H^{2}(\mathcal{D}),...,H^{n}(\mathcal{D})).$
where we denote the inclusions as $i_{i,j}$. As noted before, the freely generated Heyting algebra is simply the directed limit of this sequence. In order to generate a $\mathsf{KC}$ algebra, we need to quotient this sequence at appropriate steps. Namely, we need to replace $H^{2}(\mathcal{D})$ by the following:
$K_{C}(\mathcal{D})=H^{2}(\mathcal{D})/\{(\neg\neg i_{0,2}(a)\vee\neg i_{0,2}(a),\top):a\in\mathcal{D}\}.$
We then construct the alternative sequence:
$(\mathcal{D},H(\mathcal{D}),K_{C}(H(\mathcal{D})),K_{C}^{2}(H(\mathcal{D})),...).$
We will now dually study what this construction looks like.
Definition 43.
The motivation for this definition is the following:
Proposition 44.
Let $g:X\to Y$ be a continuous and order-preserving function, and let $\mathcal{D}=X^{*}$. Then:
$K_{C}(H(\mathcal{D}))^{*}\cong V_{g,K}(V_{g}(X))$
Proof.
By duality, we know that $K_{C}(H(\mathcal{D}))^{*}$ is obtained by considering all elements of $V_{r_{g}}(V_{g}(X))$ which satisfy the added axioms. Then we will show that for each such $C$, $C\Vdash\neg\neg[U]\vee\neg[U]$ if and only if $C$ is well-directed.
To see this, note that, formally, if $C\nVdash\neg\neg[U]\vee\neg[U]$ for some clopen upset $U\subseteq X$; this means that $C\notin[-[-U]]\cup[[-U]]$. In other words, there is some $D\in C$ such that $D\notin-[-U]$, and some $D^{\prime}\notin[-U]$. The former means that $D\subseteq-U$, and the later means that $D^{\prime}\cap U\neq\emptyset$. Let $x\in D^{\prime}\cap U$. Then $D^{\prime\prime}={\uparrow}x\cap D$ is such that $D^{\prime}\leq D^{\prime\prime}$, and so because $C$ is $r_{g}$-open, there is some $E\in C$ such that $D^{\prime}\leq E$ and $r_{g}(D^{\prime\prime})=r_{g}(E)$, i.e., $x$ is the root of $E$. Since $U$ is a clopen upset, $E\subseteq U$, and $D\subseteq-U$, so certainly
@@
Conversely, assume that $C$ is not well-directed. Let $D,D^{\prime}\in C$ be such that ${\uparrow}D\cap{\downarrow}D^{\prime}=\emptyset$. Using the properties of Priestley spaces, namely Strong Zero-Dimensionality, we can then find a clopen upset $U$ such that $D\subseteq U$ and $D^{\prime}\subseteq-U$; but this means precisely that $C\nVdash\neg\neg[U]\vee\neg[U]$ by the arguments above. ∎
Proposition 44 shows that the construction $V_{g,K}$ indeed yields a Priestley space, and so, that we can proceed inductively. In order to conclude our results, we will need to show, however, that this construction is free amongst the directed Priestley spaces. This amounts to showing the following:
Lemma 45.
Let $Z$ be a directed Priestley space, and let $h:Z\to X$ be a $g$-open, continuous and order-preserving map. Let $h^{*}:Z\to V_{g}(X)$ be the canonical map yielded by Lemma 17. Then the unique $r_{r}$-open, continuous and order-preserving map $h^{\prime}$ making the obvious triangle commute factors through $V_{g,K}(V_{g}(X))$.
@@
Similar to Lemma 17 we have that the definition of $h^{**}$ is entirely forced. All that we need to do is verify that given $x\in Z$, $h^{**}$ is well-directed.
@@
Denote by $V_{GK}(X)$ the projective limit of the complex as before. Then using these constructions we obtain:
Theorem 46.
@@
By our earlier remarks, we have that $\mathcal{K}$ is the directed union of Heyting algebras, and must satisfy the Weak Excluded Middle. To show freeness, suppose that $\mathcal{M}$ is a $\mathsf{KC}$ algebra and $k:\mathcal{D}\to\mathcal{M}$ is a map preserving the relative pseudocomplements coming from $\mathcal{A}$. Then using Lemma 45 systematically, we see that $\mathcal{M}$ will factor through $K_{C}^{n}(H(\mathcal{D}))$ for every $n$, and hence, a fortriori, through the limit. This shows the result. ∎
@@ 5.3 LC-Algebras
We finally consider the case of $\mathsf{LC}$ algebras. As is well-known, such a variety is locally finite, which means that we ought to expect our step-by-step construction to in fact stabilise at a given finite step when the original algebra is finite. As we will see, this is true in fact in greater generality. Let us say that a Priestley space $X$ is prelinear if for each $x\in X$, ${\uparrow}x$ is a linear order.
Definition 47.
@@
As before, we define
$L_{C}(\mathcal{D})=H(\mathcal{D})/\{(i_{0,1}(a)\rightarrow i_{0,1}(b)\vee i_{0,1}(b)\rightarrow i_{0,1}(a),\top):a,b\in\mathcal{D}\}$
And we cosntruct the following sequence, dual to a Vietoris-like complex:
$(\mathcal{D},L_{C}(\mathcal{D}),L_{C}^{2}(\mathcal{D}),...)$
Proposition 48.
Let $g:X\to Y$ be a continuous and order-preserving function, and let $\mathcal{D}=X^{*}$. Then:
$L_{C}(\mathcal{D})^{*}\cong V_{g,L}(X)$
@@
Assume that $U,V$ are clopen subsets of $X$, and $C\nVdash[U]\rightarrow[V]\vee[V]\rightarrow[U]$. This amounts to saying that
@@
Conversely, assume that $C$ is not linearised. Since $x\nleq y$, there is some clopen upset $U$ such that $x\in U$ and $y\notin U$, and since $y\nleq x$, there is some clopen upset $V$ such that $y\in V$ and $x\notin V$. These two subsets then witness that $C\nVdash[U]\rightarrow[V]\vee[V]\rightarrow[U]$. ∎
We now prove, again, an analogue to Lemma 17:
Lemma 49.
@@
where $V_{i}^{L}(X)$ is defined as in Definition 16 except that from the second step onwards we use $V_{g,L}(-)$ rather than $V_{g}(-)$. Then in the same way as before, we can prove:
Theorem 50.
@@
Consider a finite poset $P$. Because of local finiteness, we know that $V_{G}^{L}(P)\cong V_{n}^{L}(P)$ must hold for some finite $n$ (otherwise the construction would add infinitely many new points). By Theorem 40, $V^{L}_{1}(X)\cong X$ can never hold for such a poset, since $V_{1}^{L}(-)=V_{1}(-)$. But it is reasonable to ask whether $V_{n}^{L}(X)\cong V_{n+1}^{L}(X)$ holds for any poset. Notably, we will see that in fact the construction always stabilises in two steps.
Lemma 51.
@@
Proposition 52.
Let $X$ be a Priestley space. Assume that $V_{L}^{2}(X)$ is prelinear. Then $V_{L}^{3}(X)\cong V_{L}^{2}(X)$.
@@
We prove that the only rooted linearised and $r$-open subsets of $V_{L}^{2}(X)$ are of the form ${\uparrow}x$ for $x\in V_{L}^{2}(X)$. Let $C\subseteq{\uparrow}x$ be a linearised rooted and $r$-open subset, such that $r(C)=x$. If $C\neq{\uparrow}x$ there is some $y$ such that $x\leq y$ and $y\notin C$. Because $C$ is $r$-open, there is some $y^{\prime}\in C$ such that $r(y)=r(y^{\prime})$. Now note that $y^{\prime}$ and $y$ are comparable (they are both successors of $x$), so by Lemma 51, we have that $y=y^{\prime}$, a clear contradiction. So we must have no such $y$ must exist, and hence, that no $C$ in these conditions can exist. We then conclude that every rooted linearised subset is of the form ${\uparrow}x$ for some $x$, which yields an easy isomorphism. ∎
Proposition 53.
Let $X$ be a Priestley space. Then $V_{L}^{2}(X)$ is prelinear.
@@
Assume that $x\leq y$ and $x\leq z$, and look at $r(y)$ and $r(z)$. Since these are in $x$, which is a chain, assume that $r(y)\leq r(z)$. Then we claim that $y\leq z$. Indeed, assume that $A\in z$. Then $r(z)\leq A$, so $r(y)\leq A$. Because $y$ is $r$-open, let $A^{\prime}$ be such that $A^{\prime}\in y$, and $r(A)=r(A^{\prime})$. Then $A^{\prime}\in x$, and since $A\in z$, $A\in x$ as well; so by Lemma 51, we have that $A=A^{\prime}$. This shows that $A\in y$, which shows the result. ∎
@@
Theorem 54.
Let $X$ be a Priestley space. Then $V_{L}^{2}(X)\cong V_{L}^{\infty}(X)$.
@@
By Proposition 53 we know that $V_{L}^{2}(X)$ will be prelinear; and by Proposition 52 we have that iterating the construction past the second stage does not generate anything new. Hence for each $n>2$, $V_{L}^{n}(X)\cong V_{L}^{2}(X)$, meaning that the projective limit will be likewise isomorphic. ∎
The above theorem, coupled with some of the theory we have developed for Heyting algebras, and which could easily be generalised for Gödel algebras, can have some computational uses: given two Gödel algebras, to compute their coproduct, we can look at their dual posets, form the product, and apply $V_{L}^{2}$, to obtain the resulting product. This provides an alternative perspective from the recursive construction discussed in [DAntona2006, aguzzoligodelcoproducts], and can be compared with the work of Carai in [caraifreegodelalgebras]. In logical terms, it implies the following interesting property of Gödel logic:
Theorem 55.
Let $\varphi(p_{0},...,p_{n})$ be a formula in the language of $\mathsf{IPC}$. Then over $\mathsf{LC}$ $\varphi$ is equivalent to a formula $\psi$, over the same propositional variables, of implication degree $2$.

### v4

@@ 6.1 Boolean algebras
Our general method in handling subvarieties will be to take the Vietoris complex construction, and show that by quotienting by additional equations at appropriate levels, one can obtain sub-complexes which limit to the free algebra in the desired variety. To illustrate this, we consider the simple case of Boolean algebras: let $\mathcal{D}$ be a distributive lattice. Consider $\mathbf{F}_{DLat}(\{(a,b):a,b\in D\})/\Theta\cup\{p\vee\neg p=1\}$, the free algebra quotiented as before, with the addition of the equation $p\vee\neg p$, with which one forces Booleannness. Dually, given $X$ a Priestley space, this takes a given subspace of the space $V_{1}(X)$ (recall, by definition 3.7 that this is simply the set of closed and rooted subsets, $V_{r}(X)$):
Lemma 6.1.
Let $X$ be a Priestley space. Given a point $C\in V_{1}(X)$, we have that $C\in[\neg U]\cup[U]$ for each clopen upset $U\subseteq X$ if and only if $C$ is a singleton.
@@
Let $B_{1}(X)\subseteq V_{1}(X)$ be the set of singletons of $X$. Using Lemma 6.1 one can prove the following:
Theorem 6.2.
For each Priestley space $X$, $B_{1}(X)\cong X$ as Stone spaces, and $B_{1}(X)$ is isomorphic to the Priestley space $(X,=)$.
@@
The isomorphism is clear; the freeness follows by the same arguments as in Lemma 3.11. ∎
@@
It is also worthy to note that Boolean algebras have a peculiar feature as far as our construction $V_{G}$ goes: if $X$ is a Priestley space with a discrete order (definitionally equivalent to a Stone space), then $V_{G}(X)\cong X$, and in fact $V_{1}(X)\cong X$ already (since rooted subsets of singletons have to simply be singletons as well). Stone spaces also exhibit some further interesting properties in this respect:
Theorem 6.3.
@@
    $V_{1}(X)\cong X$ where the root map witnesses the isomorphism;
@@
Assume that $X$ is not a Stone space, i.e., there is a pair of points $x,y$ such that $x\leq y$. Note that then $x<y$ is a subposet of $X$. Recall from Example 3.1 that from such a poset one constructs the whole Rieger-Nishimura ladder, and there is throughout each layer of the construction a point $z_{n}$ such that $r$ does not provide an isomorphism. Since all sets involved are finite, this means that $V_{n}(X)$ will always likewise contain such a point $z_{n}$ where $r$ cannot provide an isomorphism.
@@ 6.2 KC-algebras
To generalise our construction to $\mathsf{KC}$ we will work with the axiomatization of this logic given by adding to $\mathsf{IPC}$ the axiom $\neg p\vee\neg\neg p$. Since this equation has rank $2$, our restriction will need to happen at the second stage, which is where one can appropriately define the law.
Definition 6.4.
The motivation for this definition is the following, where we recall that given $g:X\to Y$ a Priestley morphism, $V_{2}(X)=V_{r_{g}}(V_{g}(X))$:
Proposition 6.5.
Let $g:X\to Y$ be a continuous and order-preserving function. Given $C\in V_{2}(X)$ we have that $C\in[[X-U]]\cup[V_{1}(X)-[X-U]]$ for each clopen upset $U\subseteq X$ if and only if $C$ is well-directed.
@@
Assume that $C\notin[V_{1}(X)-[X-U]]\cup[[X-U]]$. In other words, there is some $D\in C$ such that $D\notin V_{1}(X)-[X-U]$, and some $D^{\prime}\notin[X-U]$. The former means that $D\subseteq X-U$, and the later means that $D^{\prime}\cap U\neq\emptyset$. Let $x\in D^{\prime}\cap U$. Then $D^{\prime\prime}={\uparrow}x\cap D$ is such that $D^{\prime}\leq D^{\prime\prime}$, and so because $C$ is $r_{g}$-open, there is some $E\in C$ such that $D^{\prime}\leq E$ and $r_{g}(D^{\prime\prime})=r_{g}(E)$, i.e., $x$ is the root of $E$. Since $U$ is a clopen upset, $E\subseteq U$, and $D\subseteq-U$, so certainly
@@
Conversely, assume that $C$ is not well-directed. Let $D,D^{\prime}\in C$ be such that ${\uparrow}D\cap{\downarrow}D^{\prime}=\emptyset$. Using the properties of Priestley spaces, namely Strong Zero-Dimensionality, we can then find a clopen upset $U$ such that $D\subseteq U$ and $D^{\prime}\subseteq X-U$; but this means precisely that $C\notin[[X-U]]\cup[V_{1}(X)-[X-U]]$, by the arguments above. ∎
Proposition 6.5 shows that the construction $V_{g,K}$ indeed yields a Priestley space, and so, that we can proceed inductively. In order to conclude our results, we will need to show, however, that this construction is free. Recall that we say that a Priestley space $X$ is directed if whenever $x,y,z\in X$, and $x\leq y,z$ there is some $y,z\leq w$. Then we can show:
Lemma 6.6.
Let $Z$ be a Priestley space which is directed, and let $h:Z\to X$ be a $g$-open, continuous and order-preserving map. Let $h^{*}:Z\to V_{g}(X)$ be the canonical map yielded by Lemma 3.8. Then the unique $r_{r}$-open, continuous and order-preserving map $h^{\prime}$ making the obvious triangle commute factors through $V_{g,K}(V_{g}(X))$.
@@
Similar to Lemma 3.8 we have that the definition of $h^{**}$ is entirely forced. All that we need to do is verify that given $x\in Z$, $h^{**}$ is well-directed.
@@
Denote by $V_{GK}(X)$ the projective limit of the Vietoris complex obtained by iterating the construction $V_{g,K}$, as before. Then using these constructions we obtain:
Theorem 6.7.
@@
By our earlier remarks, we have that $\mathcal{K}$ is the directed union of Heyting algebras, and must satisfy the Weak Excluded Middle. To show freeness, suppose that $\mathcal{M}$ is a $\mathsf{KC}$ algebra and $k:\mathcal{D}\to\mathcal{M}$ is a map preserving the relative pseudocomplements coming from $\mathcal{A}$. Then using Lemma 6.6 systematically, we see that $\mathcal{M}$ will factor through $K_{C}^{n}(H(\mathcal{D}))$ for every $n$, and hence, a fortriori, through the limit. This shows the result. ∎
@@ 6.3 LC-Algebras
We finally consider the case of $\mathsf{LC}$ algebras. As is well-known, such a variety is locally finite, which means that we ought to expect our step-by-step construction to in fact stabilise at a given finite step when the original algebra is finite. As we will see, much more is true. Let us say that a Priestley space $X$ is prelinear if for each $x\in X$, ${\uparrow}x$ is a linear order.
Definition 6.8.
@@
As before, we can prove:
Proposition 6.9.
Let $g:X\to Y$ be a continuous and order-preserving function. Then for each $C\in V_{1}(X)$, we have that $C\in[-U\cup V]\cup[-V\cup U]$ if and only if $C$ is linearised.
@@
Assume that $U,V$ are clopen subsets of $X$, and $C\notin[-U\cup V]\cup[-V\cup U]$. This amounts to saying that
@@
Conversely, assume that $C$ is not linearised. Since $x\nleq y$, there is some clopen upset $U$ such that $x\in U$ and $y\notin U$, and since $y\nleq x$, there is some clopen upset $V$ such that $y\in V$ and $x\notin V$. These two subsets then witness that $C\notin[-U\cup V]\cup[-V\cup U]$. ∎
We now prove, again, an analogue to Lemma 3.8:
Lemma 6.10.
@@
where $V_{i}^{L}(X)$ is defined as in Definition 3.7 except that from the second step onwards we use $V_{g,L}(-)$ rather than $V_{g}(-)$. Then in the same way as before, we can prove:
Theorem 6.11.
@@
Consider a finite poset $P$. Because of local finiteness, we know that $V_{G}^{L}(P)\cong V_{n}^{L}(P)$ must hold for some finite $n$ (otherwise the construction would add infinitely many new points). By Theorem 6.3, $V^{L}_{1}(X)\cong X$ can never hold for such a poset unless it is discrete, since $V_{1}^{L}(-)=V_{1}(-)$. But it is reasonable to ask whether $V_{n}^{L}(X)\cong V_{n+1}^{L}(X)$ holds for any poset. Notably, we will see that in fact the construction always stabilises in two steps, even for infinite spaces $X$.
Lemma 6.12.
@@
Proposition 6.13.
Let $X$ be a Priestley space. Assume that $V_{2}^{L}(X)$ is prelinear. Then $V_{3}^{L}(X)\cong V_{2}^{L}(X)$.
@@
We prove that the only rooted linearised and $r$-open subsets of $V_{2}^{L}(X)$ are of the form ${\uparrow}x$ for $x\in V_{2}^{L}(X)$. Let $C\subseteq{\uparrow}x$ be a linearised rooted and $r$-open subset, such that $r(C)=x$. If $C\neq{\uparrow}x$ there is some $y$ such that $x\leq y$ and $y\notin C$. Because $C$ is $r$-open, there is some $y^{\prime}\in C$ such that $r(y)=r(y^{\prime})$. Now note that $y^{\prime}$ and $y$ are comparable (they are both successors of $x$), so by Lemma 6.12, we have that $y=y^{\prime}$, a clear contradiction. So we must have no such $y$ must exist, and hence, that no $C$ in these conditions can exist. We then conclude that every rooted linearised subset is of the form ${\uparrow}x$ for some $x$, which yields an easy isomorphism. ∎
Proposition 6.14.
Let $X$ be a Priestley space. Then $V_{2}^{L}(X)$ is prelinear.
@@
Assume that $x\leq y$ and $x\leq z$, and look at $r(y)$ and $r(z)$. Since these are in $x$, which is a chain, assume that $r(y)\leq r(z)$. Then we claim that $y\leq z$. Indeed, assume that $A\in z$. Then $r(z)\leq A$, so $r(y)\leq A$. Because $y$ is $r$-open, let $A^{\prime}$ be such that $A^{\prime}\in y$, and $r(A)=r(A^{\prime})$. Then $A^{\prime}\in x$, and since $A\in z$, $A\in x$ as well; so by Lemma 6.12, we have that $A=A^{\prime}$. This shows that $A\in y$, which shows the result. ∎
@@
Theorem 6.15.
Let $X$ be a Priestley space. Then $V_{2}^{L}(X)\cong V_{G}^{L}(X)$.
@@
By Proposition 6.14 we know that $V_{L}^{2}(X)$ will be prelinear; and by Proposition 6.13 we have that iterating the construction past the second stage does not generate anything new. Hence for each $n>2$, $V_{n}^{L}(X)\cong V_{2}^{L}(X)$, meaning that the projective limit will be likewise isomorphic. ∎
The above theorem, coupled with some of the theory we have developed for Heyting algebras, and which could easily be generalised for Gödel algebras, can have some computational uses: given two Gödel algebras, to compute their coproduct, we can look at their dual posets, form the product, and apply $V_{L}^{2}$, to obtain the resulting product. This provides an alternative perspective from the recursive construction discussed in [DAntona2006, aguzzoligodelcoproducts], and can be compared with the work of Carai in [caraifreegodelalgebras].
In logical terms, this has some interesting consequences. Recall from [Shehtman2016] [^7] There this property is studied under the name of “finite formula depth property”; the present name was introduced since it makes obvious the connection between such a property and local tabularity, as well as allowing simpler aliases, e.g., $\mathsf{LC}$ being a $2$-uniform logic. that a superintuitionistic logic $L$ is said to be n-uniform if for each formula $\varphi$, $\varphi$ is equivalent over $L$ to a formula with implication rank $n$. We say that $L$ is uniformly locally tabular if there is some $n$ such that $L$ is $n$-uniform. It was shown by Shehtman that uniform local tabularity entails local tabularity, and it is open whether the converse holds. Our result above implies the following:
Theorem 6.16.
The logic $\mathsf{LC}$ is $2$-uniform.
This result invites the question of how one should convert an arbitrary formula into one of the normal forms of $\mathsf{LC}$. This sort of computational procedure requires a tighter understanding of the algebraic structure of the step by step construction, and we leave it for future work.